OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -3x + 4y = -19.5 -3x + y = -10.5

OpenStudy (unklerhaukus):

second equation becomes \[y=3x-10.5\] substituting this into first equation \[-3x + 4(3x-10.5) = -19.5\]\[-3x+12x-42=-19.5\] solving for x \[9x=22.5\]\[x=\frac{22.5}{9}=2.5\] substitute this into second equation \[-3(2.5) + y = -10.5\]\[y=-10.5+7.5\]\[y=-3\] Hence the solution to the system is the point \[(2.5,-3)\]

OpenStudy (radar):

The problem asked for the "elimination" method. Using the elimination method: Subtract: -3x+4y=-19.5 _ -3x + y=-10.5 ----------------------- 3y=-9 solving for y getting y=-3 You eliminated the x and solved for y. Now simply substitute the value of y (-3) into the second equation getting -3x -3 = -10.5 -3x=-10.5+3 -3x=-7.5 x=2.5